OpenStudy (anonymous):

Solve 2/(x) - 4/(x + 1) = 3

OpenStudy (anonymous):

x = 1/3 or x = -2

OpenStudy (anonymous):

how did you get that tho??

OpenStudy (anonymous):

ok ,, first( 2 ( x+1 ) - 4x ) / x ( x+1)

OpenStudy (anonymous):

do u get that ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

you take a common denominator which is x ( x+1)

OpenStudy (anonymous):

then ( 2 ( x+1 ) - 4x ) / x ( x+1) = 3

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do you get it this far ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok then expand it to be (2x + 2 - 4x ) / x^2 + x = 3

OpenStudy (anonymous):

then cross multiply we will get : 3x^2 + 3x = 2-2x

OpenStudy (anonymous):

then 3x^2 + 5x - 2 = 0 then you factorize to get (3x-1)(x+2) = 0 which will get the answers x = 1/3 or x = -2

OpenStudy (anonymous):

do you understand ? or still lost ?

OpenStudy (anonymous):

yes i do a little

OpenStudy (anonymous):

do u need more help ?

OpenStudy (anonymous):

noo i'll try the other ones by my self thx

OpenStudy (anonymous):

sure ,, glad i could help